solve the following system
6x-9y=8
14x-21y=18

Answer by Alan3354(69443)

solve the following system
6x-9y=8 --> 2x-3y = 8/3
14x-21y=18 --> 2x-3y = 18/7
-------------
Inconsistent, the graphs are parallel lines, no solution.